OpenStudy (anonymous):

olitical analysts estimate the probability that Hillary Clinton will run for president in 2008 is 45%, and the probability that NY's Governor George Pataki will run as the Republican candidate is 20%. If their political decisions are independent, then what is the probability that only Hillary runs for president?

OpenStudy (ybarrap):

Let H = Event that Hillary runs Let P = Even that Pataki runs The probability that Hillary runs is $$ H=H\cap(P\cup\bar{P})\\ \text{Therefore, }\\ =P((H\cap P)\cup (H\cap\bar P))\\ =P(H|P)P(P)+P(H|\bar P)P(\bar P)\\ =P(H)P(P)+P(H)P(\bar P)\\ =(0.45)(0.20)+(0.45)(0.80)\\ =0.45 $$ Of course. Since the probability that Hillary runs is independent of whether or not Pataki runs.

OpenStudy (anonymous):

the answer is .36

OpenStudy (ybarrap):

Okay. The probability that ONLY Hilary runs is the same as the probability that Hilary runs and Pitaki does not run $$ P(H\cap \bar P)=P(H)P(\bar P)=0.45\times0.80=0.36 $$